OpenStudy (hpaul24):

Find the 6th term of the sequence with t1 = -4 and tn = 5tn-1. -12,500 -2500 -62,500 -50,000

OpenStudy (sshayer):

\[r=\frac{ t _{n} }{ t _{n-1} }=5\] \[t _{k}=t _{1}\left( r \right)^{k-1}\] put k=6 and calculate

OpenStudy (triciaal):

@sshayer can you explain please?

OpenStudy (sshayer):

\[t _{6}=-4\left( 5 \right)^{6-1}=-4*5^5=-4 \times 3125=-12500\]